OpenStudy (anonymous):

tan(4t)-tan(7t)=1+tan(7t)tan(4t)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{tan(4t)-tan(7t)=1+tan(7t)tan(4t)\\~\\ \dfrac{tan(4t)-tan(7t)}{1+tan(7t)tan(4t)}=1\\~\\ \color{blue}{[tan(a)-tan(b)=\dfrac{tana-tanb}{1+tana+tanb}]}\\~\\ tan(4t-7t)=1\\~\\ tan(-3t)=tan(\dfrac{\pi}{4})\\~\\ tan(2\pi-3t)=tan(\dfrac{\pi}{4})\\~\\ 2\pi-3t=\dfrac{\pi}{4}\\~\\ t=\dfrac{7\pi}{12}\\~\\ }\end{align}\)